LIST OF RICHARD WIEGANDT’S PUBLICATIONS

(complete as of January 1, 2012)

Reviewed publications:

1) On complete semigroups, Acta Sci. Math. Szeged, 19 (1958), 93–97; MR20#2387, Zbl.85.252.

2) On complete semimodules, Acta Sci. Math. Szeged, 19 (1958), 219–223; MR20#6476, Zbl.85.253.

3) On the general theory of Möbius inversion formula and Möbius product, Acta Sci. Math. Szeged, 20 (1959), 164–180; MR21#7176, Zbl.87.27.

4) Bemerkung über die einstufig nichtregulären Ringe, Acta Sci. Math. Szeged, 21 (1960), 350–352; MR#A3156, Zbl.102.276.

5) Bemerkung zu einer Arbeit von Herrn Steinfeld, Acta Sci. Math. Szeged, 23 (1962), 74–75; MR26#3739, Zbl.215.380.

6) Über die Struktursätze von Halbringen, Annales Univ. Sci. Budapest., 5 (1962), 51–68; MR26#6218, Zbl.123.9.

7) Über halbeinfache linear kompakte Ringe, Studia Sci. Math. Hungar., 1 (1966), 31–38; MR34#5885, Zbl.145.267.

8) Über linear kompakte reguläre Ringe, Bull. Acad. Polon. Sci., 13 (1965), 445–446; MR32#163, Zbl.138.265.

9) Über transfinit nilpotente Ringe, Acta Math. Acad. Sci. Hungar., 17 (1966), 101–114; MR34#210, Zbl.138.264.

10) Über lokal linear kompakte Ringe, Acta Sci. Math. Szeged, 28 (1967), 255–260; MR36#3834, Zbl.153.367.

11) Vizsgálatok a lineárisan kompakt gyűrűk elméletében (Beiträge zur Theorie der linear kompakten Ringe) I, MTA III. Oszt. Közl., 16 (1966), 239–267; MR38#585, Zbl.209.342.

12) Vizsgálatok a lineárisan kompakt gyűrűk elméletében (Beiträge zur Theorie der linear kompakten Ringe) II, MTA III. Oszt. Közl., 16 (1966), 333–363; MR38#585, Zbl.209.342.

13) (with O. Steinfeld): Über die Verallgemeinerungen und Analoga der Wedderburn–Artinschen und Noetherschen Struktursätze, Math. Nachr., 34 (1967), 143–156; MR36#1495, Zbl.149.279.

14) Radical and semisimplicity in categories, Acta Math. Acad. Sci. Hungar., 19 (1968), 345–364; MR38#5882, Zbl.162.327.

15) On compact objects in categories, Publ. Math. Debrecen, 15 (1968), 267– 281; MR38#4546, Zbl.174.301.

16) (with F. Szász): On the dualization of subdirect embeddings, Acta Math. Acad. Sci. Hungar., 20 (1969), 289–302; MR40#2739, Zbl.192.339.

17) (with F. Szász): On the duality of radical and semisimple classes in categories, Acta Math. Acad. Sci. Hungar., 21 (1970), 175–182; MR41#8491, Zbl.211.31.

18) On subdirect embeddings in categories, Fund. Math., 68 (1970), 7–12; MR41#6938, Zbl.196.34.

19) A note on Kurosh–Amitsur radical properties, Annales Univ. Sci. Budapest., 12 (1969), 63–65; MR41#255, Zbl.191.34.

20) On the socle of an object in categories, Acta Sci. Math., Szeged, 31 (1970), 245–251; MR43#3326, Zbl.284.18006.

21) A note on compact objects, Publ. Math. Debrecen, 18 (1971), 99–102; MR46#5416, Zbl.249.18002.

22) Local and residual properties in bicategories, Acta Sci. Math., Szeged, 32 (1971), 195–205; MR47#6809, Zbl.216.76.

23) Structure theorems for objects, Publ. Math. Debrecen, 19 (1972), 191–198; MR50#13184, Zbl.267.18003.

24) A note on lower radicals, Annales Univ. Sci. Budapest., 13 (1970), 165–166; MR47#265, Zbl.207344.

25) On the structure of lower radical semigroups, Czechoslovak Math. J., 22 (1972), 1–6; MR45#2064, Zbl.238.2001.

26) On structure spaces, Studia Sci. Math. Hungar., 6 (1971), 205–209, MR48#5940, Zbl.237.18001.

27) Radical–semisimple classes, Periodica Math. Hungar., 3 (1973), 243–245; MR48#6159, Zbl.263.16007.

28) (with F. Szász): On hereditary radicals, Periodica Math. Hungar., 3 (1973), 235–241; MR48#6158, Zbl.263.16006; Correction: ibidem, 6 (1975), 21; MR51#5645, Zbl.301.16007.

29) Radicals coinciding with the Jacobson radical on linearly compact rings, Beiträge Alg. und Geom., Halle, 1 (1971), 195–199; MR46#1831, Zbl.228.16004.

30) Homomorphically closed semisimple classes, Studia Univ. Babeş–Bolyai, Cluj, 17 (1972), no 2, 17–20; MR49#5031, Zbl.351.16003.

31) Semisimple properties preserved by surjections, Coll. Math. Soc. J. Bolyai 6, Rings, Modules, and Radicals, Keszthely, 1971, North-Holland, 1973, 507–514; MR52#13996, Zbl.267.18002.

32) (with Eqbal Ahmed): On lower radicals of semigroups, Math. Nachr., 57 (1973), 163–167; MR48#11378, Zbl.229.20059.

33) (with M. A. Rashid): The herditariness of the upper radical, Acta Math. Acad. Sci. Hungar., 24 (1973), 343–347, MR#482173, Zbl.271.18003.

34) (with A. V. Arhangel’skiĭ): Connectednesses and disconnectednesses in topology, General Topology & Appl., 5 (1975), 9–33; MR51#4162, Zbl.329.54008.

35) • On linearly compact primitive and semisimple rings, Appendix to M. Petrich: Rings and semigroups, Springer Lecture Notes in Math., 380 (1974), 152–166; MR49#7288, Zbl.278.20068.

36) • Radical and semisimple classes of rings, Queen’s Papers in Pure & Appl. Math. 37, Kingston Ontario, 1974, pp iv+248; MR50#2227, Zbl.324.16006.

37) On N-radicals, J. Nat. Sci. & Math., Lahore, 13 (1973), 255–262; MR54#10314, Zbl.321.16007.

38) (with E. Fried): Connectednesses and disconnectednesses of graphs, Algebra Universalis, 5 (1975), 411–428; MR56#15474, Zbl.08002.

39) A general radical theory and its application in topology and graph theory, Stud. Alg. und Anwend. 1, Universale Algebren und Theorie der Radikale, Berlin, 1976, 77–84: MR56#11868, Zbl.343.08001.

40) A condition in general radical theory and its meaning for rings, topological spaces and graphs, Acta Math. Acad. Sci. Hungar., 26 (1975), 233–240; MR53#5426, Zbl.321.16008.

41) Characterization of the Baer radical by almost nilpotent rings, Publ. Math. Debrecen, 23 (1976), 15–17; MR54#2715, Zbl.34716.004; Correction: ibidem, 25 (1978), 327–328; MR80k:16014, Zbl.392.16006.

42) (with L. Márki and P. N. Stewart): Radicals and decomposability of semigroups and rings, Annales Univ. Sci. Budapest., 18 (1975), 27–36; MR55#391, Zbl.331.16007.

43) Radicals defined by means of elements, Österreich. Akad. Wiss., math.-naturw. Kl., S.-ber., Abt. II, 184 (1975), 117-125; MR55#393, Zbl.333.16008.

44) (with J. Wiesenbauer): Über die Unabhängigkeit der Distributivitätsbedingungen, Atti Accad. Naz. Rend. Cl. Fis. Mat. Natur., Roma, 58 (1975), 819–822; MR57#12354, Zbl.346.08006.

45) (with L. C. A: van Leeuwen and C. Roos): Characterizations of semisimple classes, J. Austral. Math. Soc. Ser. A, 23 (1977), 172–182; MR56#3045, Zbl.356.16003.

46) (with A. Widiger): Theory of radicals for hereditarily artinian rings, Acta Sci. Math., Szeged, 39 (1977), 305–312; MR58#16748, Zbl.347.16005.

47) (with L. Rédei and O. Steinfeld): Die wissenschaftliche Tätigkeit von Andor Kertész, Publ. Math. Debrecen, 23 (1976), 1–9; MR53#12844, Zbl.348.01015.

48) (with L. C. A. van Leeuwen): Three examples concerning the lower radical construction, Periodica Math. Hungar., 9 (1978), 121–126; MR57#3182, Zbl.328.16007.

49) (with W. G. Leavitt): Torsion theory for not necessarily associative rings, Rocky Mountain J. Math., 9 (1979), 259–271; MR80d:17005, Zbl.421.17001.

50) (with L. C. A. van Leeuwen): Constructions and compositions of radical and semisimple classes, Annales Univ. Sci. Budapest., 21 (1978), 65–75; MR80f:16014, Zbl.388.16005.

51) (with C. G. Chehata): Radical theory for fully ordered groups, Mathematica, Cluj, 20 (1978), 143–157; MR80g:06020, Zbl.409.06008.

52) (with P. N. Ánh): Linearly compact semisimple rings and regular modules, Math. Japonica, 23 (1978), 335–338; MR80d:16026, Zbl.404.16025.

53) (with T. Anderson): Weakly homomorphically closed semisimple classes, Acta Math. Acad. Sci. Hungar., 34 (1979), 329–336; MR81c:17007, Zbl.429.17002.

54) (with E. Fried): Abstract relational structures and their torsion theory, Contr. to General Algebra, J. Heyn, Klagenfurt, 1979, 97–111; MR80e:18006, Zbl.402.18007.

55) (with L. C. A. van Leeuwen): Radicals, semisimple classes and torsion theories, Acta Math. Acad. Sci. Hungar., 36 (1980), 37–47; MR82j:17004, Zbl.466.17004.

56) (with E. Fried): Abstract relational structures I, (General theory), Algebra Universalis, 15 (1982), 1–21; MR83i:18017a, Zbl.501.18002.

57) (with E. Fried): Abstract relational structures II, (Torsion theory), Algebra Universalis, 15 (1982), 22–39; MR83i18017b, Zbl.501.18003.

58) (with L. Márki): On semisimple classes of semigroups with zero, Semigroups, Springer Lecture Notes in Math., 855 (1981), 55–72; MR82g:20095, Zbl.452.20061.

59) (with T. Anderson): Semisimple classes of alternative rings, Proc. Edinburgh Math. Soc., 25 (1982), 21–26; MR84f:17011, Zbl.477.17006.

60) Semisimple classes and H-relations, Studia Sci. Math. Hungar., 13 (1978), 181–185; MR83a:16009, Zbl.423.16004.

61) (with T. Anderson): On essentially closed classes of rings, Annales Univ. Sci. Budapest., 24 (1981), 107–111; MR87m:16018, Zbl.487.16004.

62) (with L. C. A: van Leeuwen): Semisimple and torsionfree classes, Acta Math. Acad. Sci. Hungar., 38 (1981), 73–81; MR83h:17005, Zbl.481.17002.

63) (with Ju. M. Rjabuhin): On special radicals, supernilpotent radicals and weakly homomorphically closed classes, J. Austral. Math. Soc., Series A, 31 (1981), 152–162; MR82k:16009, Zbl.464.16004.

64) (with P. N. Stewart): Quasi-ideals and bi-ideals in radical theory of rings, Acta Math. Acad. Sci. Hungar., 39 (1982), 289–294; MR84i:16010, Zbl.519.16006.

65) (with B. J. Gardner): Characterizing and constructing special radicals, Acta Math. Acad. Sci. Hungar., 40 (1982), 73–83; MR84a:16009, Zbl.504.16004.

66) (with L. Márki): Remarks on radicals in categories, Category Theory, Springer Lecture Notes in Math., 962 (1982), 190–196; MR84a:18019, Zbl.502.1803.

67) (with W. Lex): Torsion theory for acts, Studia Sci. Math. Hungar., 16 (1981), 263–280; MR85g:08004a, Zbl.488.20053.

68) (with G. Betsch): Non-hereditary semisimple classes of near-rings, Studia Sci. Math. Hungar., 17 (1982), 69–75; MR85m:16020, Zbl.507.16023.

69) Near-rings and radical theory, Conf. on Near-rings and Near-fields, San Benedetto del Tronto, Italy, 1981, 49–58; MR84e:16024, Zbl.502.16028.

70) (with I. A. Amin): Torsion and torsionfree classes of acts, Contr. to General Algebra 2, B. G. Teubner, Stuttgart and Hölder-Pichler-Tempsky, Wien, 1983, 19–34; MR85g:08004a, Zbl.546.20059.

71) (with N. V. Loi): Small ideals and the Brown–McCoy radical, Coll. Math. Soc. J. Bolyai 38, Radical Theory, Eger 1982, North-Holland, 1985, 253– 264; MR88j:16012, Zbl.585.16003.

72) (with A. Suliński): Radicals of rings graded by abelian groups, Coll. Math. Soc. J. Bolyai 38, Radical Theory, Eger 1982, North-Holland, 1985, 607– 617; MR88f:16013, Zbl.586.16001.

73) (with M. M. Parmenter and P. N. Stewart): On the Groenewald–Heyman strongly prime radical, Quaest. Math., 7 (1984), 225–240; MR86:16011, Zbl.554–16005.

74) (with S. Veldsman): On the existence and non-existence of complementary radical and semisimple classes, Quaest. Math., 7 (1984), 213–224; MR86c:18004, Zbl.592.18004.

75) (with T. Anderson and K. Kaarli): Radicals and subdirect decompositions, Comm. in Algebra, 13 (1985), 479–494; MR86d:16010, Zbl.553.16004.

76) (with P. N. Ánh and N. V. Loi): On the radical theory of Andrunakievich varieties, Bull. Austral. Math. Soc., 31 (1985), 257–269; MR87e:17007, Zbl.574.17001.

77) (with B. de la Rosa): Characterizations of the Brown–McCoy radical, Acta Math. Hungar., 46 (1985), 129–132; MR87d:17002, Zbl.588.17003.

78) (with N. V. Loi): Involution algebras and the Anderson–Divinsky–Suliński property, Acta Sci. Math., Szeged, 50 (1986), 5–14; MR88a:16017, Zbl.611.17002.

79) (with P. N. Ánh): Semisimple classes of nonassociative rings and Jordan algebras, Comm. in Algebra, 13 (1985), 2669–2690; MR87c:17008, Zbl.579.17004.

80) On the commutativity of certain semigroups, Theory of Semigroups, Proc. Greifswald Conf. 1984, 125–127; MR87c:20099, Zbl.654.20069.

81) On subdirectly irreducible near-rings which are fields, Proc. Conf. on Nearrings and Near-fields, Tübingen, Germany, 1985, North-Holland, 1987, 295– 298; MR88e:160059, Zbl.615.16024.

82) On rings with subsets satisfying permutation identities, Karachi J. Math., 3 (1985), 1–7; Zbl.646.16027.

83) • (editor with L. Márki): Colloquia Mathematica Societatis János Bolyai 38, Radical Theory, Eger 1982, North-Holland, 1985; MR88e:16001, Zbl.5750009.

84) Radical theory of rings, Math. Student, 51 (1983), 145–185; MR90m:16011, Zbl.711.16010.

85) (with L. Márki and R. Mlitz): A general Kurosh–Amitsur radical theory, Comm. in Algebra, 16 (1988), 249–305; MR89h:18007, Zbl.646.08006.

86) (with L. Márki and R. Mlitz): A note on radical and semisimple classes of topological rings, Acta Sci. Math., Szeged, 51 (1987), 145–151; MR88j:16052, Zbl.645.16027.

87) (with B. de la Rosa): On dual radicals and ring elements, J. Austral. Math. Soc. Series A, 44 (1988), 164–170; MR88m:16008, Zbl.643.16003.

88) (with L. Márki): Basic concepts of Kurosh–Amitsur radicals in various structures, a historical survey, Contr. to General Algebra 4, Proc. Krems Conf. 1985, B. G. Teubner, Stuttgart and Hölder-Pichler-Tempsky, Wien, 1987, 111–124; MR89e:16017, Zbl.683.16004.

89) (with T. Anderson and K. Kaarli): On left strong radicals of nearrings, Proc. Edinburgh Math. Soc., 31 (1988), 447–456; MR89i:16034, Zbl.619.16023.

90) • Sections 42, 55, 56, 77, 78, 80 and 81 in A. Kertész: Lectures on Artinian Rings, edited by R. Wiegandt, Akadémiai Kiadó, Budapest, 1987; MR88m:16016, Zbl.681.16001.

91) Recent results in the general radical theory of rings and ring-like structures, Rings, Modules and Radicals, Proc. Hobart Conf. 1987, Pitman Res. Notes in Math. 204, Longman Sci. & Tech., 1989, 153–171; MR91b:16023, Zbl.672.16005.

92) (with K. I. Beǐdar): Splitting theorems for nonassociative rings, Publ. Math. Debrecen, 38 (1991), 121–143; MR92a:17003, Zbl.726.17003.

93) On splitting torsion theories of acts, Contr. to General Algebra 6, B. G. Teubner Stuttgart, Hölder-Pichler-Tempsky Wien, 1988, 311–316; MR91k:20080, Zbl.669.20053.

94) 0n split null extensions in radical theory, Radical Theory, Proc. 1988 Sendai Conf., Uchida Rokakuho, Tokyo 1989, 143–151; MR90d:16008, Zbl.726.16015.

95) (with N. V. Loi): On involution rings with minimum condition, Ring Theory, Israel Conf. Proc., 1 (1989), 206–214; MR91c:16036, Zbl.739.16022.

96) Kurosh–Amitsur radical theory for acts, Karachi J. Math., 5 (1989), 1–7; Zbl.739.20031.

97) (with K. I. Beǐdar): Rings with involution and chain conditions, J. Pure & Appl. Algebra, 87 (1993), 205–220; MR94g:16030, Zbl.787.16021.

98) Splitting theorems in algebra, Contr. General Algebra 7, Hölder-Pichler-Tempsky Wien, B. G. Teubner Stuttgart, 1991, 361–371; MR93b:17002, Zbl.747.16006.

99) (with B. de la Rosa and Julia van Niekerk): A concrete analysis of the radical concept, Math. Pannonica, 3(2) (1992), 3–15; Correction: ibidem, 4/1 (1993), 151; MR94i:16010, Zbl.783.16011 and 791.16018.

100) (with P. N. Ánh): Morita duality for Grothendieck categories, J. Algebra, 168 (1994), 273–293; MR95h:18011, Zbl.808.18006.

101) (with H. J. Weinert): A Kurosh–Amitsur radical theory for proper semifields, Comm. in Algebra, 20 (1992), 2419–2458; MR93h:12006, Zbl.771.16016.

102) (with H. J. Weinert): Complementary radical classes of proper semifields, Coll. Math. Soc. J. Bolyai 61, Theory of Radicals, Szekszárd 1991, North-Holland 1993, 297–310; MR94i:12005, Zbl.805.16040.

103) (with B. de la Rosa and S. Veldsman): On the theory of Plotkin radicals, Chinese J. Math., 21 (1993), 33–54; MR95c:17007, Zbl.791.16017.

104) (with Y. Fong): Subdirect irreducibility and radicals, Quaest. Math., 16 (1993), 103–113; MR94g:17002, Zbl.786.17003.

105) (with P. H. Lee): On radicals and rings with involution, PU.M.A. Ser. A, 3 (1992), 219–224; MR94i:17005, Zbl.787.17002.

106) (with Y. Fong and S. Veldsman): Radical theory in varieties of near-rings in which the constants form an ideal, Comm. in Algebra, 21 (1993), 3369–3384; MR94e:16052, Zbl.798.16033, Corrigendum: ibidem, 22 (1994), 3197–3198; MR95b:16047, Zbl.818.16033.

107) (with B. de la Rosa and Y. Fong): Complementary radicals revisited, Acta Math. Hungar., 65 (1994), 253–264; MR95f:16021, Zbl.821.16022.

108) (with Y. Fong and F. K. Huang): Radical theory for group semiautomata, Acta Cybernetica, 11 (1994), 169–188; MR95g:68076, Zbl.820.20074.

109) (with C. Roos): On the radical theory of graded rings which are inversive hemirings, Studia Sci. Math. Hungar., 32 (1996), 107–117; MR97h:16064, Zbl.854.16029.

110) • (editor with L. Márki): Colloquia Mathematica Societatis János Bolyai, 61, Theory of Radicals, Szekszárd 1991, North-Holland 1993; MR94d:16001, Zbl.778.00029.

111) On the structure of involution rings with chain condition, Tap Chí Toán Hoc (J. Math., Vietnam Math. Soc.), 21 (1993), 1–12; MR96i:16053.

112) On powers of radicals, Kyungpook Math. J., 34 (1994), 227–228; MR1325601(MathSciNet), Zbl.831.1609.

113) (with K. I. Beǐdar): ������ � ���������� � � ��������� �� ���� ���������� (Rings with involution and chain conditions on bi-ideals), ���� ���� ����, 48, 5(293), (1993), 159–160; English translation: Russian Math. Surv., 48 #5, 161–162 (1993); MR1258762(MathSciNet), Zbl.826.16031.

114) (with R. Mlitz): Semisimple classes of hypernilpotent and hyperconstant near-ring radicals, Archiv Math., 63 (1994), 414–419; MR95i:16050,

Zbl.811.16037.

115) A note on supplementing radicals, Northeast. Math. J., 11 (1995), 476– 482; MR97c:16026, Zbl.861.16014.

116) (with S. Tumurbat): On the lattice of strongly hereditary radicals, Contr. to General Algebra 9, Proc. Linz Conf. 1994, Hölder-Pichler-Tempsky Wien, B. G. Teubner Stuttgart, 1995, 309–312; MR98g:00020, Zbl.890.16012.

117) (with H. J. Weinert): On the structure of semifields and lattice-ordered groups, Periodica Math. Hungar., 32 (1996), 129–147; MR97g:12011, Zbl.896.12001.

118) (with G. F. Birkenmeier): Essential covers and complements of radicals, Bull. Austral. Math. Soc., 53 (1996), 261–266; MR97b:16026, Zbl.849.16020.

119) (with R. Mlitz and A. D. Sands): Radicals coinciding with the von Neumann regular radical on artinian rings, Monatsh. für Math., 125 (1998), 229–239; MR99d:16025, Zbl.899.16008.

120) (with H. E. Heatherly and E. K. S. Lee): Involutions on universal algebras, in: G. Saad and M. J. Thomsen (eds), Nearrings, Nearfields and K-Loops, Proc. Hamburg Conf. 1995, Kluwer 1997, 269–282; MR99c:08005, Zbl.893.08003.

121) (with K. I. Beidar): Radicals induced by the total of rings, Beiträge Alg. und Geom., 38 (1997), 149–159; MR98e:16021, Zbl.874.16014.

122) (with U. A. Aburawash): The Rees and Steinfeld theorems for semigroups with involution, Semigroup Forum, 57 (1998), 440–444; MR99g:20106, Zbl.926.20037.

123) • (editor with B. J. Gardner and Liu Shaoxue): Rings and Radicals, Proc. International Conf. Shijiazhuang ’94, Pitman Research Notes in Math. Series 346, Longman 1996; MR97b:16001, Zbl.839.00021.

124) (with G. F. Birkenmeier): Pseudocomplements in the lattice of torsion classes, Comm. in Algebra, 26 (1998), 197–220; MR98m:16036, Zbl.896.16022.

125) (with P. N. Ánh): Compactness in categories and interpretations, in: S. K. Jain and S. Tariq Rizvi (eds), Advances in Ring Theory, Birkhäuser 1997, 17–30; MR98m:18002, Zbl.887.18002.

126) (with R. Mlitz): Near-ring radicals depending only on the additive groups, Southeast Asian Bull. Math., Springer-Verlag, 22 (1998), 171–177; MR2000b:16082, Zbl.923.16038.

127) Laudatio anläßlich des 70. Geburtstages von Herrn Prof. Dr. Gy. Maurer, Grazer Math. Berichte, 334 (1998), 3–7; MR2001f:01048a, Zbl.942.01025.

128) Rings distinctive in radical theory, Quaest. Math., 22 (1999), 303–328; MR2001e:16032, Zbl.957.16016.

129) (with K. I. Beidar): Radical assignments and radical classes, Bul. Acad. Ştiinţe Rep. Moldova, Matematica, 2(30) 1999, 17–27; MR2000k:16029, Zbl.946.16013.

130) Supplementing radicals for nonassociative rings, Nonassoc. Alg. & Appl., São Paulo 1998, Lecture Notes in Pure & Appl. Math. 211, Marcel Dekker 2000, 457–469; MR2001e:17005, Zbl.1041.17003.

131) Some aspects of radical theory, XV Escola de Álgebra, Canela, Brasil 1998, Matemática Contemporânea, 16 (1999), 291–305; MR2001k:16036, Zbl.973.16015.

132) Rings with unique minimal subrings, East–West J. Math., 1 (1999), 237– 241; MR2000i:16043, Zbl.953.16022.

133) (with S. Tumurbat): Principally left hereditary and principally left strong radicals, Algebra Colloquium, 8 (2001), 409–418; MR2002h:16035, Zbl.1007.16014.

134) (with L. Márki and R. Mlitz): Brown–McCoy radicals for general near-rings, Quaest. Math., 24 (2001), 481–490; MR2002h:16071, Zbl.997.16037.

135) (with K. I. Beidar and E. R. Puczyłowski): Radicals and polynomial rings, J. Austral. Math. Soc., 72 (2002), 23–31; MR2002j:16023, Zbl.1007.16012.

136) Radical theory: developments and trends, Mathematics and the 21^{st} Century, Cairo 2000, World Scientific 2001, pp 141–152; MR2002m:16023, Zbl.1008.16022.

137) (with G. F. Birkenmeier): Supplementing radicals and decompositions of near-rings, Acta Math. Hungar., 94 (2002), 269–280; MR2003d:16054, Zbl.1017.16035.

138) (with S. Tumurbat): A note on special radicals and partitions of simple rings, Comm. in Algebra, 30 (2002), 1769–1777; MR2003a:16034, Zbl.1003.16011.

139) (with R. Mlitz): Semisimple classes containing no trivial near-rings, Studia Sci. Math. Hungar., 38 (2001), 331–337; MR1877789(MathSciNet), Zbl.1004.16048.

140) Rédei – szubjektív emlékek történelmi háttérben (Rédei – personal recollections with historical background), Mat. Lapok, 1998–99/3–4, 74–90; MR 2004 Author Index M–Z, p 2706, MR2023678(MathSciNet).

141) (with D. Isabel C. Mendes): On essential left ideals of associative rings, Math. Pannonica, 12 (2001), 217–224; MR2002j:16004, Zbl.981.16002.

142) (with S. Tumurbat): Subhereditary radicals and Brown–McCoy semisimple rings, Bul. Acad. Ştiinţe Rep. Moldova, Matematica, 3(34) 2000, 11–20; MR2003i:16034, Zbl.1025.16012.

143) • Sections C.18 (The Jacobson radical) and C.19 (General radical theory) in: A. V. Mikhalev and G. Pilz (eds), The Concise Handbook of Algebra, Kluwer Academic Publisher 2002, pp 213–219; MR2004c:00001 Zbl.1008:00004.

144) (with H. J. Weinert): A new Kurosh–Amitsur radical theory for proper semifields I, Math. Pannonica, 14 (2003), 3–28; MR2004b:16068, Zbl.1059.12003.

145) (with H. J. Weinert): A new Kurosh–Amitsur radical theory for proper semifields II, Math. Pannonica, 14 (2003), 149–164; MR2005i:16091, Zbl.1087.12003.

146) (with S. Tumurbat): On polynomial and multiplicative radicals, Quaest. Math., 26 (2003), 453–469; MR2005a:16031, Zbl.1059.1614.

147) Radical theory: main issues and recent developments, Advances in Algebra, Proc. ICM Satellite Conf., Algebra and Related Topics, Hong Kong 2002, World Scientific 2003, pp 494–505; MR2005c:00013, Zbl.1039.16017.

148) (with S. Tumurbat): On radicals with Amitsur property, Comm. in Algebra, 32 (2004), 1219–1227; MR2005g:16038, Zbl.1068.16021.

149) (with S. Tumurbat): Radicals of polynomial rings, Soochow J. Math., 29 (2003), 425–434; MR2004k:16055, Zbl.1061.16031.

150) (with S. Tumurbat): On the matric-extensibility of radicals, J. Applied Algebra & Discrete Structures, 2 (2004), 119–130; MR2005c:16038,

Zbl.1072.16019.

151) • (with B. J. Gardner): Radical theory of rings, Pure & Appl. Math. 261, Marcel Dekker 2004, pp xii+387; MR2004m:16031, Zbl.1034.16025.

152) (with S. Tumurbat): Radicals around Köthe’s Problem, Bul. Acad. Ştiinţe Rep. Moldova, Matematica, 1(44) 2004, 76–84; MR2005g:16037, Zbl.1074.16010.

153) (with Bh. Satyanarayana): On the f-prime radical of near-rings, Near-rings and Near-fields, Proc. Hamburg Conf. 2003, Springer 2005, pp 293–299; MR2006f:16064, Zbl.1082.16049.

154) (with S. Tumurbat): On A-radicals, Math. Slovaca, 56 (2006), 113–119; MR2007a:16033, Zbl.1150.16018.

155) (with K. I. Beidar, L. Márki and R. Mlitz): Primitive involution rings, Acta Math. Hungar., 109 (2005), 357–368; MR2007c:16066, Zbl.1094.16020.

156) (with S. Tumurbat): A-radicals of involution rings, Southeast Asian Bull. Math., 29 (2005), 393–399; MR2007a:16032, Zbl.1082.16028.

157) (with N. V. Loi): Subdirect irreducibility of algebras and acts with an additional unary operation, Miskolc Math. Notes, 6 (2005), 217–224; MR2006m:08004, Zbl.1095.08003.

158) (with N. V. Loi): On the Amitsur property of radicals, Algebra & Discrete Math., 3 (2006), 92–100; MR2006b:16028. Zbl.1116.16024.

159) Radical and torsion theory for acts, Semigroup Forum, 72 (2006), 312–328; MR2006k:20126, Zbl.1093.20037.

160) (with E. R. Puczyłowski): Kostia’s contribution to radical theory and related topics, Rings and Nearrings, Proc. Internat. Conf. of Algebra in Memory of Kostia Beidar, Tainan, Taiwan 2005, Walter de Gruyter, Berlin - New York 2007, pp 121–157; MR2008f:16001, Zbl.1130.16015.

161) (with L. Márki): History of the conferences on radical theory, Algebra & Discrete Math., 3 (2007), 1–10; MR2421771(MathSciNet), Zbl.1164.16001.

162) (with B. J. Gardner and J. Krempa): Open problems in radical theory, Algebra & Discrete Math., 3 (2007), 15–17; MR2421773(MathSciNet).

163) H. J. Hoehnke’s contribution to radical theory, Sci. Math. Japonicae, 68 (2008) 193–199; MR2009i:01016, Zbl.1179.20053.

Theses:

1 Vizsgálatok a lineárisan kompakt gyűrűk elméletében (Investigations in the theory of linearly compact rings), Thesis for the degree Candidate of Math. Sci., Budapest 1965.

2 Kategoria elméleti vizsgálatok (Category theoretical investigations), Thesis for the degree Doctor of Math. Sci., Budapest 1974.

Further publications, lecture notes, manuscripts, etc:

(1) Axiomatikus vizsgálatok félgyűrűkben (Axiomatische Untersuchungen in Halbringen), Preisschrift, Universität Szeged, 1954, 12 Seiten.

(2) Über die Unabhängigkeit der Distributivitätsbedingungen, Manuskript, 1954, 13 Seiten.

(3) Gondolatok a matematika-tanítás korszerűsítéséről (Thoughts about the modernization of mathematics teaching), Orosházi Táncsics Mihály Gimnázium Évkönyve, XXV. év (1962–1963), 50–54.

(4) Kertész Andor: Vorlesungen über artinsche Ringe, (book review in Hungarian), Magyar Tudomány, 13 (1968), 723–724.

(5) B. Pareigis: Kategorien und Funktoren, (book review in Hungarian), Mat. Lapok, 21 (1970), 185–186.

(6) • Representation theory of groups, Lecture Notes in Mathematics, University of Islamabad, Pakistan, 1971, pp iv + 79.

(7) Some problems in the universal algebra with ring-theoretic background, Mini-conference on Universal Algebra, Szeged, 1971, p 23.

(8) (with M. Sadiq Zia): On the terminating of the lower radical construction, unpublished manuscript, 1972, pp 6.

(9) • Lectures on radical and semisimple classes of rings, Lecture Notes in Math., University of Islamabad, Pakistan, 1972, pp 140.

(10) • Matematikai Kislexikon, (Farkas Miklós szerk., társszerzőkkel), Műszaki Kiadó, 1972; 2. kiadás: 1974; 3. kiadás: 1979.

(11) (with C. Roos): Radicals of rings defined by means of elements, manuscript, 1974, pp 17.

(12) Structure of rings and modules, manuscript noted by I. A. Assem, University of Alexandria, Egypt, 1977, pp 72.

(13) Theory of radicals and torsions, manuscript noted by I. A. Assem, University of Alexandria, Egypt, 1977, pp 41.

(14) J. S. Golan: Decomposition and dimension in module categories, (Buchbesprechung), Internationale Math. Nachr., 127 (1981), S. 21.

(15) • Lectures on radical theory, manuscript, Cairo University, Egypt, 1982, pp 35.

(16) • Radical theory of rings, Res. Report 83/1, University of Pretoria, pp 36.

(17) Radical theory, lecture delivered at the Rand Afrikaans University, Johannesburg, March 17, 1983, manuscript, pp 6.

(18) Radical theory (Revised version of aseries of talks delivered at Cairo University, Egypt, February 1982), University of Pretoria, Research Report UPWF 83/1, April 1983, pp 36.

(19) F. A. Szász: Radicals of rings, (book review), Periodica Math. Hungar., 14 (1983), 116.

(20) A matematika tankönyvek terjedelméről (Über den Umfang der mathematischen Lehrbücher), Mat. Tanítása, XXXIII (1987), 192–194.

(21) Szerkesztési feladatok algebrai úton való megoldása (Algebraic solution of geometrical construction problems), Zsuffa L. (szerk.): Rédei köri előadások 1988–1990, Tiszakécske 1990, 28–42.

(22) On the structure of involution rings with chain condition, South African Math. Soc. Notices, 22 (1990), 167–176.

(23) Az újra felfedezett ország: a Dél–Afrikai Köztársaság (The rediscovered country: the Republic of South Africa), Magyar Tudomány, 91/7, 873–882.

(24) J. S. Golan: Torsion theories, (book review), Periodica Math. Hungar., 23 (1991), 163.

(25) B. J. Gardner: Radical theory, (book review), Priodica Math. Hungar., 24 (1992), 66-67.

(26) B. J. Gardner (editor): Rings, modules and radicals, Proc. 1987 Hobart Conf., (book review) Periodica Math. Hungar., 24 (1992), 67–68.

(27) S. Kyuno (editor): Radical theory, Proc. 1988 Sendai Conf., (book review), Periodica Math. Hungar., 24 (1992), 68.

(28) Sh. Feigelstock: Additive groups of rings I and II, (book review), Periodica Math, Hungar., 24 (1992), 240–241.

(29) (with Y. Fong and F. K. Huang): Radical theory for group semiautomata, Proc. 28-th Annual Meeting, Math. Soc. R. O. C. Taiwan, Tainan, 4-th – 5-th Dec. 1992, pp 255–273.

(30) Egy kis tigris: Tajvan (A little tiger: Taiwan), Magyar Tudomány, 93/3, 343–353.

(31) Mi az algebra és mire való? (What is algebra and what is it for?), Magyar Tudomány, 93/6, 716–722.

(32) Egy 1954-es beszélgetés Szendrei Jánossal, Juhász Gyula Tanárképző Főiskola, Szeged 1995. december 18., kézirat, 4 oldal.

(33) Tudományos előadások meg az írásvetítő (Scientific talks and the overhead projector), Magyar Tudomány, 96/3, 349–350.

(34) On the structure of semigroups with involution, talk delivered on the National Conference on Algebra, Cairo 1996, manuscript, pp 8.

(35) U. Hebisch and H. J. Weinert: Halbringe – Algebraische Theorie und Anwendungen in der Informatik, (book review), Semigroup Forum, 52 (1996), 247–248; corrigiendum: ibidem 52 (1996), 397.

(36) (with K. I. Beidar): Radical assignments and radical classes, Seminarberichte Fernuniversität Hagen, 63 (1998), 45–53.

(37) Rédei László, történelmi háttérben (L. Rédei, in historical background), Természet Világa, 133/2 (2002), 61–66.

(38) Rédei László. Szubjektív emlékek történelmi háttérben, Dugonics Társaság Évkönyve, Szeged 2002, 198–210.

(39) (with H. J. Weinert), A new Kurosh–Amitsur radical theory for proper semifields I, II, Mathematik–Bericht, 2002/2, TU Clausthal.

(40) (with S. Tumurbat), Radicals of polynomial rings, The Pumplün 70 Festschrift, RWTH-Aachen Universität, Lehrstuhl A für Mathematik 2003, pp 283–291.

(41) Miért lettem matematikus? Visszaemlékezések (szerk.: Róka Sándor), Typotex 2003, 265–271.

(42) Laudatio anläßlich des 61. Geburtstages von Herrn Prof. Dr. R. Mlitz, TU Wien, 3. Juni 2005, Manuskript, 4 Seiten.

(43) Erinnerung an Hans-Jürgen Hoehnke, Halle, 25. Januar 2008, Manuskript, 2 Seiten.

(44) Science – a természettudományok integrált oktatása (Science – integrierter Unterricht der Naturwissenschaften), Természet Világa, 139/12 (2008), 558– 559; Bajai honpolgár XX:5 (2009) 15–16 (másodközlés).

(45) Képletek a fizika tanításában (Formeln im Physikunterricht), Természet Világa, 140/5 (2009), 230–231.

(46) A tribute to Professor Andrunachievici, Academicianul Vladimir Andrunachievici, Acad. Ştiinţe Moldovei, Chişinǎu 2009, pp 186–207.

(47) Hogyan lettem matematikus? kézirat, 2009, 12 oldal.

(48) Meine Erinnerungen an Andor Kertész, Georg Cantor Sonderheft, 2010, Martin-Luther-Univesität, Halle-Wittenberg. 16–23.

(49) Emlékeim Kertész Andorról, KLTE Debrecen, 2009. május 19, kézirat, 5 oldal.

(50) RSA-kriptorendszer – nyílvános kulcsú titkosírás, kézirat, 2009, 7 oldal.

(51) An application of probability theory, manuscript, pp 2.

(52) Voller Skandal, Manuskript, Lösung durch Anwendung der verallgemeinerten Möbiusschen Umkehrungformel in Acta Sci. Math. Szeged, 20 (1959), 164-180.

(53) Laudatio anläßlich der Emeritierung von Herrn Dr. R. Mlitz, TU Wien, 13. Oktober, 2010, Manuskript, 1 Seite.

(54) Kazár József, www.tancsics-ohaza.sulinet.hu/tanar/, 3 oldal.

(55) Zatykó Sándor, www.tancsics-ohaza.sulinet.hu/tanar/, 3-4 oldal; kibövített kézirat 2011, 3 oldal.

(56) A skála matematikája – das wohltemperierte Klavier, kézirat, 2011, 1 oldal.

(57) Aranymetszés, kézirat 2011, 3 oldal.

Problems (elementary, advanced and research) and solutions:

(1) A 110. feladat megoldása, Mat. Lapok, 12 (1961), 225.

(2) A 124. feladat megoldása, Mat. Lapok, 6 (1963), 166-167. 1

(3) A 137. feladat, Mat. Lapok, 14 (1963), 164, (Fried Ervinnel közösen).

(4) A 1295. feladat, Köz. Mat. Lapok, 28 (1964), 29.

(5) Advanced problem 5571, Amer. Math. Monthly, 75 (1968), 304 and ibidem 76 (1969), 203.

(6) Three problems, Coll. Math. Soc. J. Bolyai 6, Rings, Modules and Radicals, North-Holland 1973, 518–519.

(7) A 196. feladat, Mat. Lapok, 24 (1973), 151; megoldása ibidem, 25 (1974), 371–373.

(8) Two problems in the List of Problems, Kolloquium über Algebra, June 23, 1978, TU Wien.

(9) 2. feladat az 1983. évi Schweitzer Miklós Matematikai Emlékversenyen (Problem 2 on the 1983 M. Schweitzer Contest in Higher Mathematics), Mat. Lapok, 32 (1981–84), p 162.

(10) Problem in: Rings, Modules and Radicals, Proc. Hobart Conf. 1987, Longman, 1989, p 194.

(11) 1. feladat és megoldása az 1989. évi Schweitzer Miklós Matematikai Emlékversenyen (Problem 1 and its solution on the 1989 M. Schweitzer Contest in Higher Mathematics), Mat. Lapok, 1991/4, pp 21–24.

(12) 3. feladat a 2001. évi Schweitzer Miklós Matematikai Emlékversenyen (Problem 3 on the 2001 M. Schweitzer Contest in Higher Mathematics), Mat. Lapok. 2000–2001/2, pp 57–58.

(13) Problem 1, Open Problems, General Radicals, Radical Page, http://140.116.21.154/radicals, (with S. Tumurbat).

Abstracts of conference talks;

Reviews in Mathematical Reviews and Zentralblatt für Mathematik;

Interviews in newspapers and radio;

Photo: Pálma Alexandriában, Természet Világa, 116/10, 1985, front page;

CD-ROM: Mi az algebra és mire való? A tudomány egésze, válogatta és összeállította Zsolnai József, Nemzeti Tankönyvkiadó 2005.